The eucal and euscript packages

Frank Mittelbach  Rainer Schöpf  Michael Downes

Version 2.2d, 2001/10/01

1 Introduction

This package was written originally by Frank Mittelbach and Rainer Schöpf; later it was moved into the AMSFonts distribution with only minor modifications. It can be used with L^{A}TEX 2ε with no dependency on the amsmath package.

This file sets up some font shape definitions to use the Euler script symbols in math mode. These fonts are part of the AMSFonts collection which can be found on many TEX servers. It is also directly available from the AMS and from TEX user groups.

\EuScript The expected normal use of the Euler Script alphabet is as a substitute for the Computer Modern calligraphic alphabet found in cmsy. Therefore we change the meaning of \mathcal.

 \[ \mathcal{A} = \EuScript{A} \neq \CMcal{A} \]

will produce

𝒜 = 𝒜 ≠ A

Here is a complete table of the beautiful letters drawn by Hermann Zapf:

𝒜 ℬ 𝒞 𝒟 ℰ ℱ 𝒢 ℋ ℐ
𝒥 𝒦 ℒ ℳ 𝒩 𝒪 𝒫 𝒬 ℛ
𝒮 𝒯 𝒰 𝒱 𝒲 𝒳 𝒴 𝒵

2 The Implementation

Package identification.

  \NeedsTeXFormat{LaTeX2e}% LaTeX 2.09 can't be used (nor non-LaTeX)
  [1994/12/01]% LaTeX date must be December 1994 or later
  ⟨euscript⟩\ProvidesPackage{euscript}[2001/10/01 v2.2d Euler Script fonts]
  ⟨eucal⟩\ProvidesPackage{eucal}[2001/10/01 v2.2d Euler Script fonts]

We have three things to do: 1) identify the current package, 2) enlarge the font shape tables and 3) define the ⟨math alphabet identifier⟩.

The font shapes for the Euler Script medium and bold are defined in the amsfonts.fdd file which comes with the AMS font package. We repeat their default definition here for reference only.

 \DeclareFontFamily{U}{eus}{\skewchar\font'60}
 \DeclareFontShape{U}{eus}{m}{n}{%

     <5><6><7><8><9>gen*eusm%
    <10><10.95><12><14.4><17.28><20.74><24.88>eusm10}{}
 \DeclareFontShape{U}{eus}{b}{n}{%
     <5><6><7><8><9>gen*eusb%
    <10><10.95><12><14.4><17.28><20.74><24.88>eusb10}{}

\EuScript Now we define the ⟨math alphabet identifier⟩ \EuScript both for the normal and the bold math version

  \DeclareMathAlphabet\EuScript{U}{eus}{m}{n}
  \SetMathAlphabet\EuScript{bold}{U}{eus}{b}{n}

For flexibility and backward compatibility with versions 2.1c and earlier, we save the old meaning of \mathcal as \CMcal, and use \EuScript as the initial name of the new math alphabet. Notice that we don’t do any checking to make sure the previous version of \mathcal actually refers to cmsy.

  \newcommand{\CMcal}{}
  \let\CMcal=\mathcal

The psamsfonts option is intended to mean that the Y&Y/Blue Sky Research PostScript versions of the AMSFonts should be used. In that font set, the only .tfm files provided below size 10 are 5 and 7, rather than 5,6,7,8,9. See amsfonts.dtx for more discussion.

  \DeclareOption{psamsfonts}{%
    \DeclareFontFamily{U}{eus}{\skewchar\font'60}%
    \DeclareFontShape{U}{eus}{m}{n}{<-6>eusm5<6-8>eusm7<8->eusm10}{}%
    \DeclareFontShape{U}{eus}{b}{n}{<-6>eusb5<6-8>eusb7<8->eusb10}{}%
  }

Here is a table describing the action of the eucal, euscript, and eufrak packages.

Package Option  Commands provided

      eucal  none  \mathcal
      eucal     [mathcal]  \mathcal
      eucal     [mathscr]  \mathscr (\mathcal unchanged)
      euscript none  \EuScript (obsolete)
      euscript  [mathcal]  \mathcal

      eufrak  none  \mathfrak (also obsolete \EuFrak for

compatibility)

  \DeclareOption{mathcal}{\renewcommand{\mathcal}{\EuScript}}
  \DeclareOption{mathscr}{%
    \providecommand{\mathscr}{}\renewcommand{\mathscr}{\EuScript}%
  ⟨eucal⟩ \let\mathcal=\CMcal
  }

Process the package options.

  ⟨eucal⟩\ExecuteOptions{mathcal}
  \ProcessOptions

2. THE IMPLEMENTATION 3 The usual \endinput to ensure that random garbage at the end of the file

doesn’t get copied by docstrip.

  \endinput